OpenStudy (anonymous):

Prove the difference between the squares of any 2 consecutive even numbers is twice the sum of the 2 even numbers?

OpenStudy (across):

\[(n+2)^2-n^2=n^2+4n+4-n^2=2(2n+2)\]

OpenStudy (anonymous):

took me a while, but I finally understand what how you proved it. Thanks :)